# Should We Fine-Tune or RAG? Evaluating Different Techniques to Adapt LLMs for Dialogue

### v2

@@ 1 Introduction
However, recent studies have shown the shortcomings of LLMs as dialogue model surrogates as they are prone to generate toxic, biased, and irrelevant responses (Zhang et al., 2020; Mousavi et al., 2022, 2023; Lin and Chen, 2023). To adapt LLMs to dialogue types, different techniques have been employed such as in-context learning (Brown et al., 2020; Chen et al., 2023; Meade et al., 2023) and fine-tuning (Wang et al., 2022; Komeili et al., 2022; Huang et al., 2023). Furthermore, strategies such as grounding (Gopalakrishnan et al., 2019; Zhao et al., 2023) and Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Borgeaud et al., 2022) have been proposed to improve the generation quality.
@@
Adaptation of Llama2C and MistralI using fine-tuning and in-context learning in four different dialogue types and corresponding corpora;
@@ 3.1 Datasets
Task-Oriented Dialogue (TOD) We select the dataset proposed for the ninth Dialogue System Technology Challenge (Kim et al., 2020). The dataset spans over 7 domains and contains 9k multi-domain dialogues. The dialogues include turns where the system needs to access an unstructured knowledge base of 2900 documents (FAQs) to provide a correct response.
@@ 3.2 Techniques
We evaluate in-context learning and fine-tuning as techniques to adapt LLMs for response generation in the selected dialogue types. In-context learning is a technique that uses instructions and examples to condition the generation. Instead, fine-tuning further trains the model (completely or partially) on the task of interest using a smaller-scale dataset than the pre-training phase. In a dialogue setting, fine-tuning should teach the model the notion of the dialogue and the roles of the participants.
@@ 3.3 Knowledge
For the retrieved knowledge scenario, we use the Retrieval Augmented Generation (RAG) strategy. We use an off-the-shelf retriever111https://github.com/langchain-ai/langchain (model details in §A.2) to retrieve documents from the unstructured knowledge base. First, we encode all the documents considering their content together with their topic (KGD), place or service name (TOD), or title (QA) (Karpukhin et al., 2020). Then, at each turn, we retrieve the k most similar documents based on L2 distance with the encoded context. Finally, we feed the retrieved documents to the base models together with the context to generate a response.
@@ 4 Evaluation
We conduct a comparative study on the impact of in-context learning and fine-tuning to adapt LLMs for dialogues. We select Llama2C and MistralI as base LLMs and experiment in four dialogue types: ODDs, KGDs, TODs, QA. For each dialogue type, we study the impact of external knowledge, both retrieved and gold. Further details about the implementation and the resources used are available in the appendix (§A.2).
4.1 Automatic Evaluation
Table 2 reports the perplexity of Llama2C and MistralI on the test set of each dialogue type. In all dialogue types, fine-tuned models have obtained better performance compared to in-context learning. When considering the impact of external knowledge, models fine-tuned on TODs show that knowledge slightly increases perplexity. The high perplexity obtained by in-context learning models on QA can be explained because of two reasons: first, besides the knowledge, only the question is used as context; second, while the ground truths are particularly short (4.26 tokens on average), these models generate long responses, making them unlikely to include the correct answer in the first few tokens. This does not happen for fine-tuned models since they are trained to generate shorter responses. Nevertheless, the best results have been obtained by the models using gold knowledge. We report automatic evaluation results including retriever accuracy, overlap between knowledge and response tokens, and other automatic metrics in §A.3.
@@ 4.1.1 Explainability Study
To understand the contribution of each segment of the input vector (i.e. instruction, context, knowledge, topic, and dialogue state), we compute integrated gradients (Sarti et al., 2023)222We use Inseq to compute integrated gradients. of input elements and select the most contributing input tokens (top-25%). Table 4.1 reports the percentage of most contributing tokens that fall in each segment (normalized by the length of the segment). In general, in both KGD and TOD, the dialogue history is the least contributing segment, which might indicate that only a part of the history is significant for response generation. On the other hand, in KGD the topic has a higher score than the dialogue history, suggesting its importance for response generation for this dialogue type. Interestingly, MistralI gives considerably more importance to the topic than Llama2C, decreasing the importance of the knowledge segment. For the TOD type, the most contributing segment is often the knowledge, reaching over 50% with fine-tuning. This suggests that knowledge is more relevant for TOD and that relevance changes with respect to the dialogue type.
@@ 4.2 Human Evaluation
Considering the uninterpretability of automatic evaluations, we conducted a human evaluation of the generated responses to gain more insight into the models’ performance. To evaluate the responses, we use the protocol proposed by Mousavi et al. (2022), considering three of their dimensions:
Contextualization: the response includes explicit or implicit references to the context of the dialogue;
@@
According to these dimensions, we evaluate the responses for all techniques, models, and knowledge scenarios, in all dialogue types. The only exception is QA, where we do not evaluate Appropriateness since the dimension considers coherence with respect to a dialogue history but QA only has question-answer exchanges. Instead, we extend the protocol for QA by proposing a new dimension:
@@
The dimensions can either have a positive (Contextualized, Appropriate, Correct, Valid), negative (Not Contextualized, Not Appropriate, Not Correct, Not Valid) or neutral (I don’t know) answer value. For Contextualization and Appropriateness, we also ask the annotators to motivate the negative judgments with the explanations proposed in the original protocol.
We recruited 75 annotators on the Prolific platform333https://www.prolific.com/, and we assigned 5 dialogues to each annotator. After performing quality control, we approved 65 annotators and paid them 9.00£/hour (marked as good on the Prolific platform). Due to the large number of responses, each annotator evaluated a different set of model responses for a given dialogue. However, each annotator always evaluated the ground truth response for all the assigned dialogues, as a point of reference and quality control. The inter-annotator agreement measured with Fleiss’ (Fleiss, 1971) was 0.65 (substantial agreement).
As results of the human evaluation (Table 4), we report the percentage of positively judged responses (Contextualized, Appropriate, Valid) for Llama2C and MistralI when considering different adaptation techniques (Fine-Tuning and In-Context Learning) and knowledge (No Knowledge, Retrieved Knowledge, and Gold Knowledge) across different dialogue types. As for ODDs, we report no results for the Retrieved and Gold Knowledge scenarios since no knowledge was used for this dialogue type. Additional results on Correctness are reported in §A.4.
@@
Task-Oriented Dialogue (TOD) When adapting Llama2C and MistralI to TOD, the results clearly show that fine-tuning is preferable over in-context learning. In particular, if we consider the best model for each technique, when fine-tuned Llama2C generates 20% more contextualized responses, while MistralI generates 15% more. Although fine-tuned models benefit from external knowledge, Retrieved and Gold Knowledge visibly reduce contextualization of in-context learning models (at most by 30% for Llama2C and 15% for MistralI). Similar behavior can be observed for in-context learning in terms of appropriateness, where Gold Knowledge reduces Llama2C results by 15% and MistralI by 10%. This is in line with the explainability study (Table 4.1), where models adapted with in-context learning have a lower contribution from the knowledge segment than their fine-tuned version. In general, if we consider the best models for each technique, fine-tuned models generate 25% more appropriate responses.
@@ 4.3 Explaining Negative Human Judgments
Generic: the response is generic or does not contain any reference to the context (implicit or explicit);
Hallucinated: the response is inconsistent with the information contained in the context.
@@
Unhelpful: the response candidate is not helpful to fulfil the user’s request.
@@ 5 Conclusion
Due to the limited computational resources, we could experiment with 7B models, hampering us in validating our findings on larger models. Furthermore, the human evaluation results also strongly depend on the set of hired annotators. Therefore, the reproducibility of the reported results is subject to the variability of the selection of crowd workers.
@@ A.2 Implementation and resources
Models and parameters We fine-tuned the models using LoRA (rank 32 and alpha 64) for a maximum of 10 epochs with an early stopping patience of 2. We chose AdamW (Loshchilov and Hutter, 2017) as the optimizer and used a learning rate of for Llama2C and for MistralI (selected based on the performance on the development sets). To obtain an encoding for both documents and queries, we used all-mpnet-base-v2444https://www.sbert.net/docs/pretrained_models.html. We have then stored the encoded documents in a FAISS vector store (used for retrieval).
@@ A.3 Additional Automatic Evaluation
To automatically evaluate the quality of the generated text, we have considered BLEU-4 Papineni et al. (2002), F1 (i.e. unigram overlap), and ROUGE-L Lin (2004). Furthermore, we have used KF1 Shuster et al. (2021) to measure the overlap between the prediction and the knowledge selected by the annotators. For reproducibility purposes, we have computed ROUGE-L using the official implementation555https://github.com/google-research/google-research/tree/master/rouge and all the remaining metrics using ParlAI666https://parl.ai. No pre-processing was performed on the model-generated answers.
@@ A.4 Human Evaluation
Table 8 reports the results for the Correctness dimension of Human Evaluations. Except for ODD, fine-tuning tends to improve Correctness.

### v3

@@ 1 Introduction
However, recent studies have shown the shortcomings of LLMs as dialogue model surrogates as they are prone to generate toxic, biased, and irrelevant responses (Zhang et al., 2020; Mousavi et al., 2022, 2023; Lin and Chen, 2023). To adapt LLMs to dialogue types, different techniques have been employed such as in-context learning (Brown et al., 2020; Chen et al., 2023; Meade et al., 2023; Hudeček and Dusek, 2023) and fine-tuning (Wang et al., 2022; Komeili et al., 2022; Huang et al., 2023). Furthermore, strategies such as grounding (Gopalakrishnan et al., 2019; Zhao et al., 2023) and Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Borgeaud et al., 2022) have been proposed to improve the generation quality.
@@
Adaptation of Llama2C and MistralI using fine-tuning and in-context learning111The code is available at https://github.com/sislab-unitn/Fine-Tune-or-Rag in four different dialogue types and corresponding corpora;
@@ 3.1 Datasets
Task-Oriented Dialogue (TOD) We select the dataset proposed for the first track of the ninth Dialogue System Technology Challenge (Kim et al., 2020), an augmented version of MultiWOZ 2.1 (Eric et al., 2020). The dataset spans over 7 domains and contains 9k multi-domain dialogues. The dialogues include turns where the system needs to access an unstructured knowledge base of 2900 documents (FAQs) to provide a correct response.
@@ 3.2 Techniques
We evaluate in-context learning and fine-tuning as techniques to adapt LLMs for response generation in the selected dialogue types. In-context learning is a technique that uses instructions and examples to condition the generation. Instead, fine-tuning further trains the model (completely or partially) on the task of interest using a smaller-scale dataset than the pre-training phase. In a dialogue setting, fine-tuning should teach LLMs to behave as dialogue models and account for each state of the conversation between speakers.
@@ 3.3 Knowledge
For the retrieved knowledge scenario, we use the Retrieval Augmented Generation (RAG) strategy. We use an off-the-shelf retriever222https://github.com/langchain-ai/langchain (model details in §A.2) to retrieve documents from the unstructured knowledge base. First, we encode all the documents considering their content together with their topic (KGD), place or service name (TOD), or title (QA) (Karpukhin et al., 2020). Then, at each turn, we retrieve the k most similar documents based on L2 distance with the encoded context. Finally, we feed the retrieved documents to the base models together with the context to generate a response.
@@ 4 Evaluation
We conduct a comparative study on the impact of in-context learning and fine-tuning to adapt LLMs for dialogues. We select Llama2C and MistralI as base LLMs and experiment in four dialogue types: ODDs, KGDs, TODs, and QA. For each dialogue type, we study the impact of external knowledge, both retrieved and gold. Further details about the implementation and the resources used are available in the appendix (§A.2).
@@
4.1 Automatic Evaluation
Currently available automatic metrics used for the task of response generation are not interpretable and correlate poorly with human judgments (Liu et al., 2016; Sai et al., 2022; Mousavi et al., 2022). Therefore, we focus on perplexity as it is derived from the objective function used to fine-tune the models, and present other metrics in §A.3.
Table 2 reports the perplexity of Llama2C and MistralI on the test set of each dialogue type. In all dialogue types, fine-tuned models have obtained better performance compared to in-context learning. When considering the impact of external knowledge, models fine-tuned on TODs show that knowledge slightly increases perplexity. The high perplexity obtained by in-context learning models on QA can be explained by two reasons: first, besides the knowledge, only the question is used as context; second, while the ground truths are particularly short (4.26 tokens on average), these models generate long responses, making them unlikely to include the correct answer in the first few tokens. This does not happen for fine-tuned models since they are trained to generate shorter responses. Nevertheless, the best results have been obtained with gold knowledge. We report automatic evaluation results including retriever accuracy, overlap between knowledge and response tokens, and other automatic metrics in §A.3.
@@ 4.1.1 Explainability Study
To understand the contribution of each segment of the input vector (i.e. instruction, context, knowledge, topic, and dialogue state), we compute integrated gradients (Sarti et al., 2023)333We use Inseq to compute integrated gradients. of input elements and select the most contributing input tokens (top-25%). Table 4 reports the percentage of most contributing tokens that fall in each segment (normalized by the length of the segment). In general, in both KGD and TOD, the dialogue history is the least contributing segment, which might indicate that only a part of the history is significant for response generation. On the other hand, in KGD the topic has a higher score than the dialogue history, suggesting its importance for response generation for this dialogue type. Interestingly, MistralI gives considerably more importance to the topic than Llama2C, decreasing the importance of the knowledge segment. For the TOD type, the most contributing segment is often the knowledge, reaching over 50% with fine-tuning. This suggests that knowledge is more relevant for TOD and that relevance changes with respect to the dialogue type.
@@ 4.2 Human Evaluation
Considering the uninterpretability of automatic evaluations, we conducted a human evaluation of the generated responses to gain more insight into the models’ performance. Mousavi et al. (2022) proposed four dimensions to evaluate response generation based on the most common errors and qualities. We evaluate the responses using their protocol and three of their dimensions:
Contextualization: the response includes explicit or implicit references to the dialogue history (ODD, KGD, TOD) or the gold knowledge (QA);
@@
According to these dimensions, we evaluate the responses for all techniques, models, and knowledge scenarios, in all dialogue types. The only exception is QA, where we do not evaluate "Appropriateness" since the dimension considers coherence with respect to a dialogue history but QA only has question-answer exchanges. Instead, we extend the protocol444The extended protocol is available at https://github.com/sislab-unitn/Human-Evaluation-Protocol/tree/v1.1 by proposing a new dimension for QA:
@@
For TOD we do not include a dimension to evaluate whether the response is in line with user requirements, as this can be measured automatically (via dialogue state tracking metrics e.g., Joint Goal Accuracy). The dimensions can either have a positive or negative answer value, as well as "I don’t know" to avoid forcing erroneous judgments on any of the two sides. For "Contextualization" and "Appropriateness", we also ask the annotators to motivate the negative judgments with the explanations proposed in the original protocol. We present the explanations and related results in §4.3.
We recruited 75 annotators on the Prolific platform555https://www.prolific.com/, and we assigned 5 dialogues to each annotator. After performing quality control, we approved 65 annotators with a compensation of 9.00£/hour (marked as good on the Prolific platform). Due to the large number of responses, each annotator evaluated a different set of model responses for a given dialogue. For the purpose of quality control, for each dialogue type, two dialogues were overlapping among five annotators, while the remaining dialogues were annotated by one crowd-worker with an overlap only on the ground truth. The inter-annotator agreement measured with Fleiss’ (Fleiss, 1971) was 0.65 (substantial agreement).
As results of the human evaluation (Table 4), we report the percentage of positively judged responses (Contextualized, Appropriate, Valid) for Llama2C and MistralI when considering different adaptation techniques (Fine-Tuning and In-Context Learning) and knowledge (No Knowledge, Retrieved Knowledge, and Gold Knowledge) across different dialogue types. As for ODDs, we report no results for the Retrieved and Gold Knowledge scenarios since no knowledge was used for this dialogue type. Additional results on "Correctness" are reported in §A.4.
@@
Task-Oriented Dialogue (TOD) When adapting Llama2C and MistralI to TOD, the results clearly show that fine-tuning is preferable over in-context learning. In particular, if we consider the best model for each technique, when fine-tuned Llama2C generates 20% more contextualized responses, while MistralI generates 15% more. Although fine-tuned models benefit from external knowledge, Retrieved and Gold Knowledge visibly reduce contextualization of in-context learning models (at most by 30% for Llama2C and 15% for MistralI). Similar behavior can be observed for in-context learning in terms of appropriateness, where Gold Knowledge reduces Llama2C results by 15% and MistralI by 10%. This is in line with the explainability study (Table 4), where models adapted with in-context learning have a lower contribution from the knowledge segment than their fine-tuned version. In general, if we consider the best models for each technique, fine-tuned models generate 25% more appropriate responses.
@@
These findings indicate that the best technique depends on the dialogue type and the base LLM. Regarding the techniques, in-context learning leads to more contextualized and appropriate responses in ODDs, while fine-tuning improves contextualization and appropriateness in TODs. Regarding the base LLMs, in KGDs adapting Llama2C with in-context learning leads to the best results, while MistralI benefits the most from fine-tuning. Furthermore, in QA the quality of knowledge impacts contextualization and validity the most, while adaptation techniques have a minor effect.
@@ 4.3 Explaining Negative Human Judgments
Generic: the response is generic or does not contain any reference (implicit or explicit) to the dialogue history (ODD, KGD, TOD) or the gold knowledge (QA);
Hallucinated: the response is inconsistent with the information contained in the dialogue history (ODD, KGD, TOD) or the gold knowledge (QA).
@@
Unhelpful: the response candidate is not helpful in fulfilling the user’s request.
@@ 5 Conclusion
Due to the limited computational resources, we could experiment with 7B models, hampering us in validating our findings on larger models. Furthermore, the reproducibility of human evaluation results may be subject to variability, due to possible differences in the set of crowd workers.
Acknowledgments
We acknowledge the support of the MUR PNRR project FAIR - Future AI Research (PE00000013) funded by the NextGenerationEU.
@@ A.2 Implementation and resources
Models and parameters We fine-tuned the models using LoRA (rank 32 and alpha 64) for a maximum of 10 epochs with an early stopping patience of 2. We chose AdamW (Loshchilov and Hutter, 2017) as the optimizer and used a learning rate of for Llama2C and for MistralI (selected based on the performance on the development sets). To obtain an encoding for both documents and queries, we used all-mpnet-base-v2666https://www.sbert.net/docs/pretrained_models.html. We have then stored the encoded documents in a FAISS vector store (used for retrieval).
@@ A.3 Additional Automatic Evaluation
To automatically evaluate the quality of the generated text, we have considered BLEU-4 Papineni et al. (2002), F1 (i.e. unigram overlap), and ROUGE-L Lin (2004). Furthermore, we have used KF1 Shuster et al. (2021) to measure the overlap between the prediction and the knowledge selected by the annotators. For reproducibility purposes, we have computed ROUGE-L using the official implementation777https://github.com/google-research/google-research/tree/master/rouge and all the remaining metrics using ParlAI888https://parl.ai. No pre-processing was performed on the model-generated answers.
@@ A.4 Human Evaluation
Table 8 reports the results for the "Correctness" dimension of Human Evaluations. Except for ODD, fine-tuning tends to improve correctness.
Table 9 presents the question and the answer options for the proposed "Validity" dimension used in QA.
@@
| Dimension | Question | Answer Option | Option Definition |
| Validity | Is the response candidate valid? | Valid | The response candidate includes the right information from the context to adequately answer the proposed question. |
| Not Valid | The response candidate does not include the right information from the context to adequately answer the proposed question. | ||
| I don’t know | The response candidate includes some information that is adequate to answer the proposed question, but some that is not. |